# Generalizations of Koga’s version of the Wiener-Ikehara theorem

### v2

@@ 7. Further extensions and concluding remarks
Theorem 1.1 is also covered by (B.2). In fact, suppose that (1.4) and (1.6) are satisfied. Let $\varphi$ be a real-valued even non-negative smooth function with support on $(-\lambda,\lambda)$ such that $\varphi(t)=1$ for $t\in[-\lambda/2,\lambda/2]$. Let $\widehat{\varphi}(x)=\int_{-\infty}^{\infty}\varphi(t)e^{-itx}\mathrm{d}x$, so that $\widehat{\varphi}$ is a Schwartz function. Then, since $U(\sigma+it)-g(t)\geq 0$ in the considered range,
@@
Similarly, (b.2) is always implied by (1.13) and
$\sum_{n=1}^{\infty}\frac{|c_{n}|}{n^{k}}<\infty\qquad\mbox{for some }k\in\mathbb{N}.$
@@
Koga originally stated his Tauberian theorem (cf. [22, Theorem 2]) by only imposing the boundary requirements for the Laplace transform on a sequence tending to $1^{+}$. More precisely, in addition to (1.3), he assumes [^6] His formulation of (7.1) in [22, Theorem 2] is slightly different, but equivalent in view of the well known completeness of the $L^{1}$-spaces. the existence of $\sigma_{n}\to 1^{+}$ such that
@@
We have however that (7.1) is equivalent to (1.5) in our case. In fact, since (1.3), or more generally (1.6) which allows us to view $e^{-x}S(x)$ as a tempered distribution, ensures [35, Section 6.6.9, p. 100] that $\mathcal{L}\{S;s\}$ has distributional boundary values on $\Re e\>s>1$, the relation (1.5) might be inferred from (7.1) by using a standard localization argument together with the (distributional) Schwarz reflection principle [31] (see also [8, 27] for generalized reflection principles).
Although (7.2) appears to be weaker than (1.4), Koga’s original set of hypotheses are also covered by Theorem 1.2. To see his, the same argument employed in Remark 7.1 clearly yields, say,
$\sup_{n\in\mathbb{N}}\int_{-2\lambda/3}^{2\lambda/3}|U(\sigma_{n}+it)|\mathrm{d}t<\infty,$
which in turn (via localization and the reflection principle once more) implies that (B.2) holds with $\lambda$ replaced by, say, $\lambda/2$.
@@
As Koga, we could also have only assumed in Theorem 1.2 and Theorem 1.3 that (1.4) just holds on a sequence $\sigma=\sigma_{n}\to 1^{+}$. Exactly the same argument given in Section 4 would then still yield that the boundary distribution of $U$ is a Radon measure. This comment also applies to Theorem 1.5.
Remark 7.4.
In this article we have chosen to work within the framework of distributional boundary values because distribution theory should be well-known to most analysts. We could also have worked with ultradistributions, which allows us to relax the assumption (1.7) in (B.3) to
(7.3) $U(\sigma+it)=O\left(e^{\eta\left(\frac{1}{\sigma-1}\right)}\right),\qquad\mbox{for }t\in(-\lambda,\lambda)\mbox{ and }\sigma\in(1,2],$
where $\eta:[0,\infty)\to[1,\infty)$ is non-decreasing and satisfies the mild integral growth condition
(7.4) $\int_{1}^{\infty}\frac{\log\eta(x)}{x^{2}}\mathrm{d}x<\infty.$
Likewise, Theorem 1.5 also holds if we replace (1.12) in the hypotheses (b.3) by the weaker growth condition
(7.5) $U(re^{i\theta})=O\left(e^{\eta\left(\frac{1}{1-r}\right)}\right),\qquad\mbox{for }\theta\in(-\theta_{0},\theta_{0})\mbox{ and }r\in[0,1).$
In particular, (7.3) holds with $\eta(y)=1+\sup_{x>0}(\omega(x)-x/y)$ if $\omega:[0,\infty)\to[0,\infty)$ is non-decreasing and concave with
(7.6) $\int_{1}^{\infty}\frac{\omega(x)}{x^{2}}\mathrm{d}x<\infty$
and if
(7.7) $\int_{1}^{\infty}{|S(x)|}e^{-x-\omega(x)}\mathrm{d}x<\infty.$
(To see that (7.4) holds, one can apply [7, Theorems 1.8 and 1.10].) Similarly, (7.5) holds with this $\eta$ if
(7.8) $\sum_{n=0}^{\infty}|c_{n}|e^{-\omega(n)}<\infty.$
For example, $\omega(x)=x^{\alpha/(1+\alpha)}$ with $\alpha>0$ gives rise to $\eta(y)\asymp y^{\alpha}.$
We refer to [8, 36] for the theory of boundary values of harmonic functions in spaces of ultradistributions [^7] To be able to apply those results under (7.4), one needs to pass through the majorization results [7, Theorem 1.10, p. 92], [4, Lemma 8.5, p. 233], and [8, Lemma 2.4, p. 664]..
Finally, we mention that the comments from Remark 7.1 also apply to (7.7) and (7.8), due to the non-quasianalytic condition (7.6), which, by Beurling’s theorem [2], guarantees the existence of cut-off functions $\varphi\in\mathcal{D}(\mathbb{R})$ such that $\widehat{\varphi}(x)=O(e^{-\omega(x)}).$
@@
from relatively mild regularity boundary properties of the Laplace transform. For instance, such criteria play an important role in abstract analytic number theory (see e.g. [13, Chapter 11] for applications to Beurling primes). The following extension of [10, Proposition 3.1] (cf. [13, Theorem 10.1]) could then be useful in that respect. We call $S$ log-linearly boundedly decreasing if there is $h>0$ such that
@@
That the conditions are necessary is easy to verify. Let us show their sufficiency for (7.9). Set $\widehat{\varphi}(x)=\int_{-\infty}^{\infty}\varphi(t)e^{-itx}\mathrm{d}x$ and $\Delta(x)=e^{-x}S(x)$. As shown inside the proof of [10, Proposition 3.1], $\Delta(x)=O(1)$ as $x\to\infty$ would follow from the log-linear bounded decrease if we show that $(\Delta\ast\widehat{\varphi})(h)=O(1)$ as $h\to\infty$. Also, it is shown there that we may assume without any loss of generality that $\Delta\geq 0$. Thus (the use of Parseval’s relation is justified by [10, Lemma 3.4])
@@
For example, Proposition 7.6 could have been used in Section 4 to directly show that $U(s)$ has distributional boundary values under the hypothesis (B.2) without having to pass through the conformal map argument, because once $T(x)=e^{-x}S(x)=O(1)$, the Laplace transform $\mathcal{L}\{S;s\}$ tends to the (distributional) Fourier transform of $T$ on $\Re e\>s=1$.

### v3

@@ 7. Concluding remarks
Theorem 1.1 is also directly covered by (B.2). In fact, suppose that (1.4) and (1.3) are satisfied. Let $\varphi$ be a real-valued even non-negative smooth function with support on $(-\lambda,\lambda)$ such that $\varphi(t)=1$ for $t\in[-\lambda/2,\lambda/2]$. Let $\widehat{\varphi}(x)=\int_{-\infty}^{\infty}\varphi(t)e^{-itx}\mathrm{d}x$, so that $\widehat{\varphi}$ is a Schwartz function. Then, since $U(\sigma+it)-g(t)\geq 0$ in the considered range,
@@
Koga originally stated his Tauberian theorem (cf. [17, Theorem 2]) by only imposing the boundary requirements for the Laplace transform on a sequence tending to $1^{+}$. More precisely, in addition to (1.3), he assumes [^5] His formulation of (7.1) in [17, Theorem 2] is slightly different, but equivalent in view of the well known completeness of the $L^{1}$-spaces. the existence of $\sigma_{n}\to 1^{+}$ such that
@@
We have however that (7.1) is equivalent to (1.5) in our case. In fact, since (1.3), which allows us to view $e^{-x}S(x)$ as a tempered distribution, ensures [31, Section 6.6.9, p. 100] that $\mathcal{L}\{S;s\}$ has distributional boundary values on $\Re e\>s>1$, the relation (1.5) might be inferred from (7.1) by using a standard localization argument together with the (distributional) Schwarz reflection principle [27] (see also [5, 23] for generalized reflection principles). Likewise, the condition (7.2) follows from (1.4) and (1.3) (where one might need to subtract a constant from $g$ resulting from the application of the reflection principle). Alternatively, one can also see that Koga’s original set of hypotheses is covered by Theorem 1.2 via the same argument employed in Remark 7.1, which clearly yields, say,
$\sup_{n\in\mathbb{N}}\int_{-2\lambda/3}^{2\lambda/3}|U(\sigma_{n}+it)|\mathrm{d}t<\infty.$
The latter condition in turn implies (via localization and the reflection principle once more) that (B.2) holds with $\lambda$ replaced by, say, $\lambda/2$.
@@
As Koga, we could also have only assumed in Theorem 1.2 and Theorem 1.3 that (1.4) just holds on a sequence $\sigma=\sigma_{n}\to 1^{+}$. Exactly the same argument given in Section 4 would then still yield that the boundary distribution of $U$ is a Radon measure. This comment also applies to Corollary 1.4 and Theorem 1.5.
@@
from relatively mild regularity boundary properties of the Laplace transform. For instance, such criteria play an important role in abstract analytic number theory (see e.g. [8, Chapter 11] for applications to Beurling primes). The following extension of [6, Proposition 3.1] (cf. [8, Theorem 10.1]) could then be useful in that respect. We call $S$ log-linearly boundedly decreasing if there is $h>0$ such that
@@
That the conditions are necessary is easy to verify. Let us show their sufficiency for (7.3). Set $\widehat{\varphi}(x)=\int_{-\infty}^{\infty}\varphi(t)e^{-itx}\mathrm{d}x$ and $\Delta(x)=e^{-x}S(x)$. As shown inside the proof of [6, Proposition 3.1], $\Delta(x)=O(1)$ as $x\to\infty$ would follow from the log-linear bounded decrease if we show that $(\Delta\ast\widehat{\varphi})(h)=O(1)$ as $h\to\infty$. Also, it is shown there that we may assume without any loss of generality that $\Delta\geq 0$. Thus (the use of Parseval’s relation is justified by [6, Lemma 3.4])
@@
For example, Proposition 7.5 could have been used in Section 4 to directly show that $U(s)$ has distributional boundary values under (B.2) without having to pass through the conformal map argument, because once $T(x)=e^{-x}S(x)=O(1)$, the Laplace transform $\mathcal{L}\{S;s\}$ tends to the (distributional) Fourier transform of $T$ on $\Re e\>s=1$.